OpenStudy (anonymous):

The graph of a quadratic function results in a: a. Hyperbola b. Parabola c. Ellipse d. Conical Ellipse

OpenStudy (solomonzelman):

quadratic equation, you mean. It is a vertical parabola, opening up or down.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y=x^2+4x+7}\) for example. \(\normalsize\color{blue}{ y=x^2+4x+4+3}\) \(\normalsize\color{blue}{ y=(x+2)^2+3}\) centered at (-2,3)

OpenStudy (solomonzelman):

Hyperbola and Ellipse aren't functions. (that also)